# Learning Theory for Kernel Bilevel Optimization

### v1

@@ C.1 Boundedness and Lipschitz continuity of $h^{\star}_{\omega}$ and $\hat{h}_{\omega}$
Boundedness of $\left\|h^{\star}_{\omega}\right\|_{\mathcal{H}}$ and $\left\|\hat{h}_{\omega}\right\|_{\mathcal{H}}$. Let $\omega\in\text{hull}(\Omega)$. Using Lemma G.2, we know, for any $h\in\mathcal{H}$, that:
@@
Using the expression of the partial derivative $\partial_{h}L_{in}$ established in Proposition A.1, we obtain:
@@
By Assumption (A), $K$ is bounded by $\kappa$. Hence, Jensen’s inequality yields:
@@
By Assumption (B), $\Omega$ and $\mathcal{Y}$ are compact, which implies that $\text{hull}(\Omega)\times\mathcal{Y}$ is compact. From Assumption (C), we know that the function $(\omega,y)\mapsto\partial_{v}\ell_{in}(\omega,0,y)$ is continuous. Given that every continuous function on a compact space is bounded, we obtain:
@@
To prove that $\left\|\hat{h}_{\omega}\right\|_{\mathcal{H}}\leq\frac{B\sqrt{\kappa}}{\lambda}$, we follow a similar approach to that of $\left\|h^{\star}_{\omega}\right\|_{\mathcal{H}}\leq\frac{B\sqrt{\kappa}}{\lambda}$. More precisely, we investigate the case where the expectation is with respect to the empirical estimate ^$\hat{\mathbb{P}}_{n}$ of $\mathbb{P}$.
$h^{\star}_{\omega}(x)$ and $\hat{h}_{\omega}(x)$ belong to $\mathcal{V}$. Let $\omega\in\text{hull}(\Omega)$ and $x\in\mathcal{X}$. By the reproducing property, the Cauchy-Schwarz inequality, and Assumption (A), we have:
@@
Proposition C.2 (Lipschitz continuity of $\omega\mapsto h^{\star}_{\omega}$).
Under Assumptions (B), (D), (A) and (C), the function $\omega\mapsto h^{\star}_{\omega}$ is $\frac{L\sqrt{\kappa}}{\lambda}$-Lipschitz continuous on $\textnormal{hull}(\Omega)$, where $L\coloneqq\sup_{\omega\in\textnormal{hull}(\Omega),v\in\mathcal{V},y\in\mathcal{Y}}\left\|\partial_{\omega,v}^{2}\ell_{in}(\omega,v,y)\right\|>0$, and $\mathcal{V}$ is the compact interval introduced in Proposition C.1.
@@
Let $\omega\in\text{hull}(\Omega)$. Using Propositions A.1 and A.3, we know that $h\mapsto L_{in}(\omega,h)$ is $\lambda$-strongly convex and Fréchet differentiable. Also, by Proposition A.2, $\partial_{h}L_{in}$ is Fréchet differentiable on $\mathbb{R}^{d}\times\mathcal{H}$, and, a fortiori, Hadamard differentiable. Then, by the functional implicit differentiation theorem (Petrulionytė et al., 2024, Theorem 2.1), the Jacobian $\partial_{\omega}h^{\star}_{\omega}:\mathcal{H}\to\mathbb{R}^{d}$ can be expressed as:
@@
where the first line uses the sub-multiplicative property of the operator norm $\|\cdot\|_{\operatorname{op}}$, the second line stems from the fact that $h\mapsto L_{in}(\omega,h)$ is $\lambda$-strongly convex, for any $\omega\in\mathbb{R}^{d}$, as proved in Proposition A.3, the third line follows from Proposition A.2, the fourth line uses Jensen’s inequality, and the last line is a direct consequence of the boundedness of $K$ by $\kappa$ (Assumption (A)). According to Proposition C.1, $h^{\star}_{\omega}(x)\in\mathcal{V}\coloneqq\left[-\frac{B\kappa}{\lambda},\frac{B\kappa}{\lambda}\right]$, which is a compact interval of $\mathbb{R}$, where $B\coloneqq\sup_{\omega\in\text{hull}(\Omega),y\in\mathcal{Y}}\left|\partial_{v}\ell_{in}(\omega,0,y)\right|>0$. By Assumption (B), $\Omega$ and $\mathcal{Y}$ are compact sets, hence $\text{hull}(\Omega)\times\mathcal{V}\times\mathcal{Y}$ is compact. Besides, by Assumption (C), $(\omega,v,y)\mapsto\partial_{v}\ell_{in}(\omega,v,y)$ is continuous over the domain $\text{hull}(\Omega)\times\mathcal{V}\times\mathcal{Y}$. Since every continuous function on a compact set is bounded, this leads to:
@@
Substituting this bound into Equation 14 means that $\frac{L\sqrt{\kappa}}{\lambda}$ is an upper-bound on $\left\|\partial_{\omega}h^{\star}_{\omega}\right\|_{\operatorname{op}}$. Thus, the result follows as desired. ∎
C.2 Local boundedness and Lipschitz properties of $\ell_{in}$, $\ell_{out}$, and their derivatives
Proposition C.3 (Local boundedness).
Under Assumptions (B), (D), (A) and (C), the functions $(\omega,x,y)\mapsto\ell_{out}(\omega,h_{\omega}^{\star}(x),y)$, $(\omega,x,y)\mapsto\partial_{\omega}\ell_{out}(\omega,h_{\omega}^{\star}(x),y)$, and $(\omega,x,y)\mapsto\partial_{v}\ell_{out}(\omega,h^{\star}_{\omega}(x),y)$ are bounded over $\textnormal{hull}(\Omega)\times\mathcal{X}\times\mathcal{Y}$ by some positive constant $M_{out}$. Similarly, the functions $(\omega,x,y)\mapsto\partial_{v}\ell_{in}(\omega,h^{\star}_{\omega}(x),y)$, $(\omega,x,y)\mapsto\partial_{v}^{2}\ell_{in}(\omega,h^{\star}_{\omega}(x),y)$, and $(\omega,x,y)\mapsto\partial_{\omega,v}^{2}\ell_{in}(\omega,h^{\star}_{\omega}(x),y)$ are bounded over $\textnormal{hull}(\Omega)\times\mathcal{X}\times\mathcal{Y}$ by some positive constant $M_{in}$. The constants $M_{out}$ and $M_{in}$ are defined as:
@@ C.2 Local boundedness and Lipschitz properties of $\ell_{in}$, $\ell_{out}$, and their derivatives
where $\mathcal{V}\subset\mathbb{R}$ is the compact interval defined in Proposition C.1.
@@
By Proposition C.1, we have that $h^{\star}_{\omega}(x)\in\mathcal{V}\coloneqq\left[-\frac{B\kappa}{\lambda},\frac{B\kappa}{\lambda}\right]\subset\mathbb{R}$, for any $x\in\mathcal{X}$. From Assumption (C), we know that $\ell_{in}$, $\ell_{out}$, and their partial derivatives are all continuous on $\text{hull}(\Omega)\times\mathcal{V}\times\mathcal{Y}$. Also, $\mathcal{Y}$ is compact by Assumption (B). Thus, $\text{hull}(\Omega)\times\mathcal{V}\times\mathcal{Y}$ is compact. As every continuous function defined over a compact space is bounded, we obtain that:
$\sup_{\omega\in\text{hull}(\Omega),x\in\mathcal{X},y\in\mathcal{Y}}{\left\lvert\ell_{out}(\omega,h_{\omega}^{\star}(x),y)\right\rvert}$ $\leq\sup_{\omega\in\text{hull}(\Omega),v\in\mathcal{V},y\in\mathcal{Y}}{\left\lvert\ell_{out}(\omega,v,y)\right\rvert}<+\infty,$
@@
Proposition C.4 (Local Lipschitz continuity).
Under Assumptions (B), (D), (A) and (C), there exists a positive constant $\operatorname{Lip}_{out}$ so that for any $(x,y)$ in $\mathcal{X}\times\mathcal{Y}$, the functions $\omega\mapsto\ell_{out}(\omega,h_{\omega}^{\star}(x),y)$, $\omega\mapsto\partial_{\omega}\ell_{out}(\omega,h_{\omega}^{\star}(x),y)$, and $\omega\mapsto\partial_{v}\ell_{out}(\omega,h^{\star}_{\omega}(x),y)$ are locally $\frac{\operatorname{Lip}_{out}}{\lambda}$-Lipschitz continuous over $\textnormal{hull}(\Omega)$. Similarly, there exists a positive constant $\operatorname{Lip}_{in}$ so that for any $(x,y)$ in $\mathcal{X}\times\mathcal{Y}$, the functions $\omega\mapsto\partial_{v}\ell_{in}(\omega,h^{\star}_{\omega}(x),y)$, $\omega\mapsto\partial_{v}^{2}\ell_{in}(\omega,h^{\star}_{\omega}(x),y)$, and $\omega\mapsto\partial_{\omega,v}^{2}\ell_{in}(\omega,h^{\star}_{\omega}(x),y)$ are locally $\frac{\operatorname{Lip}_{in}}{\lambda}$-Lipschitz continuous $\textnormal{hull}(\Omega)$. The constants $\operatorname{Lip}_{out}$ and $\operatorname{Lip}_{in}$ are defined, for any $0<\lambda\leq\Lambda$, as:
@@
with $M_{in}$ and $M_{out}$ being the positive constants defined in Proposition C.3, and $\mathcal{V}\subset\mathbb{R}$ is the compact interval defined in Proposition C.1.
@@
where the first line uses the chain rule, the second line applies the triangle inequality and the reproducing property of the RKHS $\mathcal{H}$, the third line follows from Proposition C.3 to bound the derivatives of $\ell_{out}$, from Proposition C.2, which states that the function $\omega\mapsto h^{\star}_{\omega}$ is $\frac{L\sqrt{\kappa}}{\lambda}$-Lipschitz continuous with $L\coloneqq\sup_{\omega\in\text{hull}(\Omega),v\in\mathcal{V},y\in\mathcal{Y}}\left\|\partial_{\omega,v}^{2}\ell_{in}(\omega,v,y)\right\|<M_{in}$, to bound ${\left\lVert\partial_{\omega}h^{\star}_{\omega}\right\rVert}_{\operatorname{op}}$, and from Assumption (A) to bound ${\left\lVert K(x,\cdot)\right\rVert}_{\mathcal{H}}$, and the last line is a direct consequence of $0<\lambda\leq\Lambda$. In a similar way, we obtain:
${\left\lVert\nabla_{\omega}\partial_{\omega}\ell_{out}(\omega,h^{\star}_{\omega}(x),y)\right\rVert}_{\operatorname{op}}\leq\frac{\bar{M}_{out}\left(\Lambda+M_{in}\kappa\right)}{\lambda},\quad{\left\lVert\nabla_{\omega}\partial_{v}\ell_{out}(\omega,h^{\star}_{\omega}(x),y)\right\rVert}\leq\frac{\bar{M}_{out}\left(\Lambda+M_{in}\kappa\right)}{\lambda},$
${\left\lVert\nabla_{\omega}\partial_{v}\ell_{in}(\omega,h^{\star}_{\omega}(x),y)\right\rVert}\leq\frac{M_{in}\left(\Lambda+M_{in}\kappa\right)}{\lambda},\quad{\left\lVert\nabla_{\omega}\partial_{v}^{2}\ell_{in}(\omega,h^{\star}_{\omega}(x),y)\right\rVert}\leq\frac{\bar{M}_{in}\left(\Lambda+M_{in}\kappa\right)}{\lambda},$
@@
Appendix D Generalization Properties

### v3

@@ D.1 Boundedness and Lipschitz continuity of $h^{\star}_{\omega}$ and $\hat{h}_{\omega}$
Boundedness of $\left\|h^{\star}_{\omega}\right\|_{\mathcal{H}}$ and $\left\|\hat{h}_{\omega}\right\|_{\mathcal{H}}$. Let $\omega\in\text{hull}(\Omega)$. Using Lemma˜H.2, we know, for any $h\in\mathcal{H}$, that:
@@
Using the expression of the partial derivative $\partial_{h}L_{in}$ established in Proposition˜B.1, we obtain:
@@
By Assumption˜(B), $K$ is bounded by $\kappa$. Hence, Jensen’s inequality yields:
@@
By Assumption˜(C), $\mathcal{Y}$ is compact, which implies that $\text{hull}(\Omega)\times\mathcal{Y}$ is compact. From Assumption˜(D), we know that the function $(\omega,y)\mapsto\partial_{v}\ell_{in}(\omega,0,y)$ is continuous. Given that every continuous function on a compact space is bounded, we obtain:
@@
To prove that $\left\|\hat{h}_{\omega}\right\|_{\mathcal{H}}\leq\frac{B\sqrt{\kappa}}{\lambda}$, we follow a similar approach to that of $\left\|h^{\star}_{\omega}\right\|_{\mathcal{H}}\leq\frac{B\sqrt{\kappa}}{\lambda}$. More precisely, we investigate the case where the expectation is with respect to the empirical estimate $\widehat{\mathbb{P}}_{n}$ of $\mathbb{P}$.
$h^{\star}_{\omega}(x)$ and $\hat{h}_{\omega}(x)$ belong to $\mathcal{V}$.
Let $\omega\in\text{hull}(\Omega)$ and $x\in\mathcal{X}$. By the reproducing property, the Cauchy-Schwarz inequality, and Assumption˜(B), we have:
@@
Proposition D.2 (Lipschitz continuity of $\omega\mapsto h^{\star}_{\omega}$).
Under Assumptions˜(A), (C), (E), (B) and (D), the function $\omega\mapsto h^{\star}_{\omega}$ is $\frac{L\sqrt{\kappa}}{\lambda}$-Lipschitz continuous on $\textnormal{hull}(\Omega)$, where $L\coloneqq\sup_{\omega\in\textnormal{hull}(\Omega),v\in\mathcal{V},y\in\mathcal{Y}}\left\|\partial_{\omega,v}^{2}\ell_{in}(\omega,v,y)\right\|>0$, and $\mathcal{V}$ is the compact interval introduced in Proposition˜D.1.
@@
Let $\omega\in\text{hull}(\Omega)$. Using Propositions˜B.1 and B.3, we know that $h\mapsto L_{in}(\omega,h)$ is $\lambda$-strongly convex and Fréchet differentiable. Also, by Proposition˜B.2, $\partial_{h}L_{in}$ is Fréchet differentiable on $\mathbb{R}^{d}\times\mathcal{H}$, and, a fortiori, Hadamard differentiable. Then, by the functional implicit differentiation theorem [ioffe1979theory, petrulionyte2024functional, Theorem 2.1], the Jacobian $\partial_{\omega}h^{\star}_{\omega}:\mathcal{H}\to\mathbb{R}^{d}$ can be expressed as:
@@
where the first line uses the sub-multiplicative property of the operator norm $\|\cdot\|_{\operatorname{op}}$, the second line stems from the fact that $h\mapsto L_{in}(\omega,h)$ is $\lambda$-strongly convex, for any $\omega\in\mathbb{R}^{d}$, as proved in Proposition˜B.3, the third line follows from Proposition˜B.2, the fourth line uses Jensen’s inequality, and the last line is a direct consequence of the boundedness of $K$ by $\kappa$ (Assumption˜(B)). According to Proposition˜D.1, $h^{\star}_{\omega}(x)\in\mathcal{V}\coloneqq\left[-\frac{B\kappa}{\lambda},\frac{B\kappa}{\lambda}\right]$, which is a compact interval of $\mathbb{R}$, where $B\coloneqq\sup_{\omega\in\text{hull}(\Omega),y\in\mathcal{Y}}\left|\partial_{v}\ell_{in}(\omega,0,y)\right|>0$. By Assumption˜(C), $\mathcal{Y}$ is a compact set, hence $\text{hull}(\Omega)\times\mathcal{V}\times\mathcal{Y}$ is compact. Besides, by Assumption˜(D), $(\omega,v,y)\mapsto\partial_{v}\ell_{in}(\omega,v,y)$ is continuous over the domain $\text{hull}(\Omega)\times\mathcal{V}\times\mathcal{Y}$. Since every continuous function on a compact set is bounded, this leads to:
@@
Substituting this bound into Equation˜13 means that $\frac{L\sqrt{\kappa}}{\lambda}$ is an upper bound on $\left\|\partial_{\omega}h^{\star}_{\omega}\right\|_{\operatorname{op}}$. Thus, the result follows as desired. ∎
D.2 Local boundedness and Lipschitz properties of $\ell_{in}$, $\ell_{out}$, and their derivatives
Proposition D.3 (Local boundedness).
Under Assumptions˜(A), (C), (E), (B) and (D), the functions $(\omega,x,y)\mapsto\ell_{out}(\omega,h_{\omega}^{\star}(x),y)$, $(\omega,x,y)\mapsto\partial_{\omega}\ell_{out}(\omega,h_{\omega}^{\star}(x),y)$, and $(\omega,x,y)\mapsto\partial_{v}\ell_{out}(\omega,h^{\star}_{\omega}(x),y)$ are bounded over $\textnormal{hull}(\Omega)\times\mathcal{X}\times\mathcal{Y}$ by some positive constant $M_{out}$. Similarly, the functions $(\omega,x,y)\mapsto\partial_{v}\ell_{in}(\omega,h^{\star}_{\omega}(x),y)$, $(\omega,x,y)\mapsto\partial_{v}^{2}\ell_{in}(\omega,h^{\star}_{\omega}(x),y)$, and $(\omega,x,y)\mapsto\partial_{\omega,v}^{2}\ell_{in}(\omega,h^{\star}_{\omega}(x),y)$ are bounded over $\textnormal{hull}(\Omega)\times\mathcal{X}\times\mathcal{Y}$ by some positive constant $M_{in}$. The constants $M_{out}$ and $M_{in}$ are defined as:
@@ D.2 Local boundedness and Lipschitz properties of $\ell_{in}$, $\ell_{out}$, and their derivatives
where $\mathcal{V}\subset\mathbb{R}$ is the compact interval defined in Proposition˜D.1.
@@
By Proposition˜D.1, we have that $h^{\star}_{\omega}(x)\in\mathcal{V}\coloneqq\left[-\frac{B\kappa}{\lambda},\frac{B\kappa}{\lambda}\right]\subset\mathbb{R}$, for any $x\in\mathcal{X}$. From Assumption˜(D), we know that $\ell_{in}$, $\ell_{out}$, and their partial derivatives are all continuous on $\text{hull}(\Omega)\times\mathcal{V}\times\mathcal{Y}$. Also, $\mathcal{Y}$ is compact by Assumption˜(C). Thus, $\text{hull}(\Omega)\times\mathcal{V}\times\mathcal{Y}$ is compact. As every continuous function defined over a compact space is bounded, we obtain that:
$\sup_{\omega\in\text{hull}(\Omega),x\in\mathcal{X},y\in\mathcal{Y}}\left\lvert\ell_{out}(\omega,h_{\omega}^{\star}(x),y)\right\rvert$ $\leq\sup_{\omega\in\text{hull}(\Omega),v\in\mathcal{V},y\in\mathcal{Y}}\left\lvert\ell_{out}(\omega,v,y)\right\rvert$$\infty$,
@@
Proposition D.4 (Local Lipschitz continuity).
Under Assumptions˜(A), (C), (E), (B) and (D), there exists a positive constant $\operatorname{Lip}_{out}$ so that for any $(x,y)$ in $\mathcal{X}\times\mathcal{Y}$, the functions $\omega\mapsto\ell_{out}(\omega,h_{\omega}^{\star}(x),y)$, $\omega\mapsto\partial_{\omega}\ell_{out}(\omega,h_{\omega}^{\star}(x),y)$, and $\omega\mapsto\partial_{v}\ell_{out}(\omega,h^{\star}_{\omega}(x),y)$ are locally $\frac{\operatorname{Lip}_{out}}{\lambda}$-Lipschitz continuous over $\textnormal{hull}(\Omega)$. Similarly, there exists a positive constant $\operatorname{Lip}_{in}$ so that for any $(x,y)$ in $\mathcal{X}\times\mathcal{Y}$, the functions $\omega\mapsto\partial_{v}\ell_{in}(\omega,h^{\star}_{\omega}(x),y)$, $\omega\mapsto\partial_{v}^{2}\ell_{in}(\omega,h^{\star}_{\omega}(x),y)$, and $\omega\mapsto\partial_{\omega,v}^{2}\ell_{in}(\omega,h^{\star}_{\omega}(x),y)$ are locally $\frac{\operatorname{Lip}_{in}}{\lambda}$-Lipschitz continuous over $\textnormal{hull}(\Omega)$. The constants $\operatorname{Lip}_{out}$ and $\operatorname{Lip}_{in}$ are defined, for any $0<\lambda\leq\Lambda$, as:
@@
with $M_{out}$ and $M_{in}$ being the positive constants defined in Proposition˜D.3, and $\mathcal{V}\subset\mathbb{R}$ is the compact interval defined in Proposition˜D.1.
@@
where the first line uses the chain rule, the second line applies the triangle inequality and the reproducing property of the RKHS $\mathcal{H}$, the third line follows from Proposition˜D.3 to bound the derivatives of $\ell_{out}$, from Proposition˜D.2, which states that the function $\omega\mapsto h^{\star}_{\omega}$ is $\frac{L\sqrt{\kappa}}{\lambda}$-Lipschitz continuous with $L\coloneqq\sup_{\omega\in\text{hull}(\Omega),v\in\mathcal{V},y\in\mathcal{Y}}\left\|\partial_{\omega,v}^{2}\ell_{in}(\omega,v,y)\right\|<M_{in}$, to bound $\left\lVert\partial_{\omega}h^{\star}_{\omega}\right\rVert_{\operatorname{op}}$, and from Assumption˜(B) to bound $\left\lVert K(x,\cdot)\right\rVert_{\mathcal{H}}$, and the last line is a direct consequence of $0<\lambda\leq\Lambda$. In a similar way, we obtain:
$\left\lVert\nabla_{\omega}\partial_{\omega}\ell_{out}(\omega,h^{\star}_{\omega}(x),y)\right\rVert_{\operatorname{op}}\leq\frac{\bar{M}_{out}\left(\Lambda+M_{in}\kappa\right)}{\lambda},\left\lVert\nabla_{\omega}\partial_{v}\ell_{out}(\omega,h^{\star}_{\omega}(x),y)\right\rVert\leq\frac{\bar{M}_{out}\left(\Lambda+M_{in}\kappa\right)}{\lambda},$
$\left\lVert\nabla_{\omega}\partial_{v}\ell_{in}(\omega,h^{\star}_{\omega}(x),y)\right\rVert\leq\frac{M_{in}\left(\Lambda+M_{in}\kappa\right)}{\lambda},\left\lVert\nabla_{\omega}\partial_{v}^{2}\ell_{in}(\omega,h^{\star}_{\omega}(x),y)\right\rVert\leq\frac{\bar{M}_{in}\left(\Lambda+M_{in}\kappa\right)}{\lambda},$
@@
Appendix E Generalization Properties